b) If the company opts for a share repurchase, how much will your total investment be worth? (5 points)
Built Rite Corp. is evaluating an extra dividend versus a share repurchase. In either case, $5,500 would be spent. Current earnings are $0.80 per share, and the stock currently selts for $33 per share. There are 250 shares outstanding. Ignore taxes and other imperfections. You own one share of stock in this company.
te Corp. is evaluating an extra dividend versus a share repurchase. In either case, $5,500 would spent. Current earnings are $0.80 per share, and the stock currently seits for $33 per share. There are 250 shares outstanding. Ignore taxes and other imperfections. You own one share of stock in this company a) If the company issues the cash dividend, how much wll your totalinvestment be worth? (5 points)
b) If the company opts for a share repurchase, how much will your total investment be worth? (5 points)
Answer #1
a) Cash received as dividend for the one share held = 5500/250 = $          22.00
Share price after dividend = (250*33-5500)/250 = $          11.00
Worth of total investment $          33.00
b) Cash received on offering the share for repurchase $          33.00